Shift Summary for Owl shift, 9/11/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Sat, 11 Sep 1999 05:48:58 -0700 (MST)

Shift Summary for Owl shift, 9/11/99

On Shift: Albert Erwin, Sydney Taegar

Hours of beam delivered at all intensities: 8.0/8.5
Hours of beam delivered at nominal intensity: 8.0
Hours of E832 data (special runs not included): 7.25
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

KTeV: 45 min - Diagnosis and access to swap ET board for CsI channel 1435.
MCR: 30 min - Problems taking us out of access.

========================================================================
Runs ended this shift:
Run 14499 - 20.3M events - BEAM_832

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

Sydney and Albert replace ET board to fix missing dynode
signal for 1435.

========================================================================
Overview of shift:

Smooth E832 running interrupted by access to swap ET board
for 1435. Channel 1435 inefficient for run 14499 spills 305 - 362.
Channel 920 inefficient for spills 360ish - EOR.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14499 00:00 BEAM_832 run in progress.
AE and previous shiftees waiting for ST to arrive.

00:06 ST arrives...

01:30 NM2D1 error in EPICURE: 3603A is below
3606A lower limit. Inform MCR - they
bring it up a few amps and then notice
that autotune was off. Beam targetting
has been ok though.

305 01:35 Beginning at spill 305, continuous
HCC inefficient alarms for CsI 1435.
1435 has no dynode signal at monitoring
station; anode signal is fine. Consult
with Ed,Julie. Cycling HV for 1435
has no effect. Decide to make access
to check signal at ET.

01:55 AE and ST on access to investigate missing
dynode signal. Problem is blown input
buffer on ET board in TE crate slot 2.
Swap ET board.

02:15 Call out of access. Notice that someone
has been tampering with the ET test stand --
the HP pulse generator is missing!

02:25 MCR calls - having trouble taking us out of
access mode.

363 02:42 Beam returns.

04:30 Discover intermittent (~every 10-20 spills)
"HCC is inefficient"
warnings since beam returned are real:
CsI 920 (large channel) has missing dynode
signal at monitoring station. ST earlier
checked but failed to notice missing HCC
bit in laser illumination on kdisp. :(
On consultation with Ed, decide to wait
for a downtime/end of shift to fix.

06:00 Mount onsplit tape QKNL34.

07:00 Tapes full. Run 14499 ends.
Call for access to swap ET board for channel 920.

========================================================================
Follow up Notes on Previous Shift

Forgot to update pedestals.

========================================================================
Items to Follow up on Next Shift

Monitor HCC hotbits and inefficient channels carefully.
High temperature and multiple power cycling during
9/10 day shift may have caused some damage.

Update pedestals after run/during downtime since we forgot
to do so in the frenzy of swapping ET boards.